# CuLEmo: Cultural Lenses on Emotion - Benchmarking LLMs for Cross-Cultural Emotion Understanding

### v1

@@ 6 Discussion
6.1 Emotion Varies Across Languages and Cultures
Data analysis from the CuLEmo dataset in Figure 2 shows notable differences in how annotators from different countries perceive the same event. For instance, Germany has the highest proportion of neutral (no emotion) labels. Figure 4 further illustrates the distribution of positive, negative, and neutral sentiments across 400 instances in each country.
How are emotions distributed across languages and cultures? Based on the dataset analysis of emotion distribution across languages, shown in Figure 2: German (87%), Arabic (58%), and English (50.5%) data have the most neutral (no emotion). Amharic (29.5%), Arabic (22%), and Spanish (21.5%) languages have the most anger emotion. These findings confirm that a single event can evoke distinct emotional reactions depending on cultural background and language.
Figure 4: Sentiment distribution across countries in the CuLEmo dataset (400 total instances).
@@ 6.2 Prompt Language Strongly Affects Cultural Emotion Understanding
As illustrated in Table 2, prompt language plays a major role in LLM performance for emotion prediction and sentiment analysis. For less-resourced languages like Hindi and Amharic, prompting in English consistently yields better results—sometimes by as much as a 20% improvement. Conversely, in-language prompts with explicit country context tend to work best for high-resource languages such as German and Spanish. These discrepancies stem from differences in both linguistic coverage and instruction-following abilities learned during pre-training. One practical solution is to leverage English prompts while specifying the target country (e.g., “You live in <<country name>>,”).
@@ 6.3 Performance Gaps Reflect Under-Represented cultures
We observe notably lower accuracy for Ethiopia and India in both emotion and sentiment tasks, suggesting that models may be less familiar with cultural practices and norms for these under-represented contexts. Ensuring greater diversity in training corpora is key to improving model performance for such cultures. Providing explicit country context can partially offset these gaps by nudging models to incorporate relevant cultural knowledge.
Overall, our findings underscore the importance of cultural context in developing and deploying LLMs. Beyond balanced data collection, researchers may explore culture-specific tuning or reinforcement learning from human feedback to further refine models’ ability to interpret and respect cultural nuances.
@@ 7 Conclusion
In this paper, we evaluated a diverse set of state-of-the-art LLMs for their ability to predict culturally aware emotion prediction and sentiment analysis tasks. We investigated the influence of including explicit country references “You live in <<country name>>” and varying the query language. Our results indicate that LLMs tend to excel at culturally driven emotions that are well-represented in their training data and underperform for less represented cultures. Specifically, we found that 1) Emotion is culture-dependent and can vary notably across languages and regions, 2) LLMs exhibit sizable performance gaps when tested on culture-specific emotions from under-represented locales, 3) Providing explicit country context in prompts improves both emotion and sentiment prediction, and 4) Sentiment analysis is better for the models, likely because it involved few class (positive, negative, neutral) than fine-grained emotion categories.
@@
Additional constraints arise from our decision to limit the emotion label space to six classes; including more categories (e.g. surprise or disgust) could yield more fine-grained insights Niu et al. (2024). Our dataset also covers only six languages/countries and comprises 400 events, which may restrict generalizability. Data bias is another concern, as emotions and cultures vary widely and cannot be exhaustively captured.
@@ Appendix A Dataset Examples
The below box shows English examples from CuLEmo datasets.
@@ B.1 Hyper-parameters
For open-source LLMs, we used the default generation hyper-parameters (top-p sampling with p = 0.9 and temperature = 0.0, max_new_tokens=200). For others, we directly employed their pre-defined interfaces, either through their online API or the CHAT function from the Transformers library. The proprietary models are called through their API in Python. All open-source models are evaluated using default parameters from the hugging face.
@@ B.2 LLM Versions
  • Gemini1.5-flash Google Deepmind (2024) — gemini-1.5-flash-002
  • GPT-4 OpenAI (2024) — gpt-4 (turbo-2024-04-09)
@@ Appendix D English Prompt Results
Figure 5: English prompting results with You live in <<country name>> context for the corresponding languages and LLMs.

### v2

@@ 6 Discussion
6.1 Variance Across Languages and Cultures
Data analysis from the CuLEmo dataset in Figure 2 shows notable differences in how annotators from different countries perceive the same event differently. For instance, the events annotated from Germany have the highest proportion of neutral (no emotion) labels. Figure 5 further illustrates the distribution of positive, negative, and neutral sentiments across 400 instances in each country.
How are emotions distributed across languages and cultures? Based on the dataset analysis of emotion distribution across languages, shown in Figure 2: German (87%), Arabic (58%), and English (50.5%) data have the most neutral (no emotion). Amharic (29.5%), Arabic (22%), and Spanish (21.5%) languages have the most anger emotion. These findings confirm that a single event can evoke distinct emotional reactions depending on the cultural background and language.
@@ 6.2 Prompt Language Strongly Affects Cultural Emotion Understanding
As illustrated in the Table 4 and summarized results in Figure 4, prompt language plays a major role in LLM performance for emotion prediction and sentiment analysis. For less-resourced languages like Amharic and Hindi, prompting in English consistently yields better results—sometimes by as much as a 20% improvement. Conversely, in-language prompts with explicit country context tend to work best for high-resource languages such as German and Spanish. These discrepancies stem from differences in both linguistic coverage and instruction-following abilities learned during pre-training. One practical solution is to leverage English prompts while specifying the target country (e.g., “You live in <<country name>>,”).
@@ 6.3 Performance Gaps Reflect Under-Represented cultures
We observe notably lower accuracy for Ethiopia and India in both emotion and sentiment tasks, suggesting that models may be less exposed to cultural practices and norms for these under-represented contexts. Ensuring greater diversity in training corpora is key to improving model performance for such cultures. Providing explicit country context can partially offset these gaps by nudging models to incorporate relevant cultural knowledge.
Overall, our findings underscore the importance of cultural context in developing and deploying LLMs. Beyond balanced data collection, researchers may explore culture-specific tuning or reinforcement learning from human feedback to further refine the abilities of models to interpret and respect cultural nuances.
@@ 7 Conclusion
In this paper, we evaluate a diverse set of state-of-the-art LLMs for their ability to predict culturally aware emotion prediction and sentiment analysis tasks. We investigate the influence of including explicit country references “You live in <<country name>>” and varying the query language. Our results indicate that LLMs tend to excel at culturally driven emotions that are well-represented in their training data and underperform for less represented cultures. Specifically, we find that 1) emotion is culture-dependent and can vary notably across languages and regions; 2) LLMs exhibit sizable performance gaps when tested on culture-specific emotions from under-represented locales; 3) providing explicit country context in prompts improves both emotion and sentiment prediction; and 4) sentiment analysis is better for the models, likely because it involved fewer class (positive, negative, neutral) than fine-grained emotion categories.
@@
Limited number of events
Our test comprises only 400 questions for each language, which is certainly not sufficient to capture the full cultural differences in emotional expression.
Drawback of majority vote
We decide the final label of the annotations using majority vote, such as an emotion label greater than or equal to three votes from a total of five annotators per instance will pass as a final emotion label. As a general drawback of the majority vote, this will exclude the perspectives of minority votes. Modeling annotator-level data without applying the majority vote can address this.
Limited emotion label space
Additional constraints arise from our decision to limit the emotion label space to six classes; including more emotion categories (e.g. surprise or disgust) during the annotation could yield more fine-grained insights Niu et al. (2024). Our dataset also covers only six languages/countries and comprises 400 events, which may restrict generalizability.
Annotation bias
Emotion annotation is subjective in nature and can vary widely depending on personal background; it likely still has consistency issues, affecting the reliability of the evaluations.
Limited model evaluations
@@
Ethics Statement
We conducted this work with careful attention to ethical considerations involving data creation, annotation, and potential downstream impacts.
1. Data Collection and Annotation
Cultural Respect
The CuLEmo dataset was curated with input from native speakers and cultural representatives. We designed questions to capture diverse cultural norms and emotional responses without perpetuating stereotypes.
Consent and Compensation
We used Amazon Mechanical Turk (MTurk) and an in-house annotation platform for data labeling. Workers were informed of the task’s nature and compensated fairly at a rate of \$12/hour, which exceeds minimum-wage standards in the majority of the annotators’ countries of residence.
Privacy and Confidentiality
All scenario-based questions were artificially created or adapted from publicly available cultural information. No personally identifiable information was collected, and no real names or private details were used.
2. Fair Representation and Potential Biases
Under-Representation
While we included six languages (English, Arabic, Amharic, German, Hindi, and Spanish) to broaden cultural coverage, bias in representation is inevitable in such datasets and evaluations. Clearly, many global cultures and languages remain unrepresented in our work, including minority language speakers of the languages we studied and speakers of those languages in less dominant regions. Additionally, cultures are complex and not subject to clean delineation. We therefore make no contention that this is a complete or fully representative dataset.
Subjectivity of Emotions
Emotions are inherently subjective and influenced by personal and cultural backgrounds. Crowd-sourced annotations may inadvertently amplify majority cultural norms or obscure minority perspectives. We minimized these risks by providing clear guidelines, but acknowledge that subjective variation is inevitable.
3. Responsible Use of the Dataset and Models
Cultural Sensitivity
The dataset includes prompts and scenarios potentially sensitive to specific cultural contexts (e.g., religious practices, social norms). We urge researchers and practitioners to exercise cultural sensitivity and caution when using the dataset or resulting models in applications that could impact cultural or ethnic groups.
Downstream Applications
Models trained or evaluated on CuLEmo could be applied in contexts such as mental health or social support, potentially affecting vulnerable populations. We encourage developers to consider safety, fairness and informed consent when deploying such systems. We caution against deployment in high-stakes settings, particularly without appropriate safeguards, user testing, and especially ethical oversight.
4. Transparency and Future Work
Open Access
We release the CuLEmo dataset publicly to facilitate reproducibility and encourage further research in culturally aware NLP.
Ongoing Improvement
Future efforts should expand cultural and linguistic diversity, refine annotation protocols, and include more nuanced emotional labels. We welcome community feedback to improve both the dataset and modeling approaches.
We aim to advance culturally aware NLP through responsible data practices, fair representation, and transparent sharing, and hope this work fosters a more inclusive understanding of emotion across languages and cultures.
Acknowledgments
This work is partially funded by the NAACL regional fund 2024. We gratefully acknowledge computing resource support provided by the Computing Research Center (CIC) at the National Polytechnic Institute (IPN) and the Marian E. Koshland Integrated Natural Sciences Center at Haverford College. We extend our sincere thanks to the Digital Scholarship team at Haverford College, particularly Anna Lacy and Patricia Guardiola, for their assistance throughout this project. We thank Professor Xerxes Minocher at Haverford College for their valuable inputs. Grissom is funded by NSF grant 2403439.
@@ Appendix A Dataset Examples
The section below shows English examples from CuLEmo datasets.
@@ B.1 Hyper-parameters
For open-source LLMs, we used the default generation hyperparameters (top-p sampling with p = 0.9 and temperature = 0.0, max_new_tokens=200). For others, we directly employed their pre-defined interfaces, either through their online API or the CHAT function from the Transformers library. The proprietary models are called through their API in Python. All open-source models are evaluated using default parameters from the hugging face.
@@ B.2 LLM Versions
  • Gemini1.5-flash Team et al. (2024) — gemini-1.5-flash-002
  • GPT-4 Achiam et al. (2024) — gpt-4 (turbo-2024-04-09)
@@ Appendix D English Prompt Results
Figure 6: English prompting results with You live in <<country name>> context for the across languages and LLMs.